OpenStudy (anonymous):

Solve Someone PLEASE HELP ME!! Ed's income from 2 stocks each year totals $280. Stock A pays dividends at the rate of 5% and stock B at the rate of 6%. If she has invested a total of $5000, how much is invested in each stock?

OpenStudy (mertsj):

x= amount in Stock A 5000-x= amount in Stock B .05x= dividends from Stock A .06(5000-x)= dividends from Stock B .05x+.06(5000-x)=280

OpenStudy (mertsj):

Solve that equation.

OpenStudy (radar):

What did you get?

OpenStudy (anonymous):

A 2000 B 3000

OpenStudy (radar):

Well they do add up to $5000.

OpenStudy (radar):

The dividends add up to $280/ Looks like a good answer. Congratulations.